OpenStudy (anonymous):

if there is a swimming pool and it contains 1 litre of water , the amount of water in the pool increases to twice the previous amount every minute , and if the pool is completely filled after 60 minutes , when does the pool reaches half the amount of water? :D

OpenStudy (harsimran_hs4):

let`s start tell me the amount of water in the pool when time is 2min = ? 3min = ?

OpenStudy (mww):

At 0 min, V = 1L At 1 min, V = 2L At 2 min, V = 4L = 2^2L At 3 min, v = 8L = 2^3L Can you see then pattern? At 60 min, what do you get for v? v(60) = 2^60:L You want to find the time for half of v(60) so this would be 2^60 / 2 = 2^59. I'll leave the last step for you. At what time in minutes would we get 2^59L of water?